The country of Cincinnatisland (country C) joins country B to form a large open economy, while country A forms a small open economy on its own with SC=50+200r and            IC=25-400r

  1. What is the new equilibrium interest rate?
  2. Is country A a borrower or a lender?
Answer #1

Saving = 50 + 200r

Investment = 25 - 400r

a) At equilibrium, saving = investment

50 + 200r = 25 - 400r

600r = -25

r = -4.17%

b) At this rate of Interest, saving = 58.33 while investment is 8.33

As saving > investment, country A is lender.